OpenStudy (anonymous):

The graph of which function has y = –1 as an asymptote?

OpenStudy (ash2326):

@monkeywall101 Are there any choices given?

OpenStudy (anonymous):

oh yea sorry (A) y=e^-x (B) y=(-x) / (1-x) (C) y=ln(x+1) (D) y= x / (x+1) (E) y= x/ (1-x)

OpenStudy (ash2326):

put x= \(\infty\) and see which one gives you y=-1

OpenStudy (anonymous):

ok thanks so much :)

OpenStudy (ash2326):

Did you get it? which is the correct option?

OpenStudy (anonymous):

im trying it out now

OpenStudy (ash2326):

okay, take your time, let me know if you need help

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

i got the last option as the answer

OpenStudy (ash2326):

Yeah, that's correct. Good work:D

OpenStudy (anonymous):

thank you! :)